During a flood a tree trunk of mass 100 kg falls down a waterfall.The waterfall is 5 m high.


If air resistance is ignored,calculate:

a.) the potential energy of the tree trunk at the top of the waterfall.

b.) the kinetic energy of the tree trunk at the bottom of the waterfall.

c.) the magnitude of the velocity of the tree trunk at the bottom of the waterfall.


1
Expert's answer
2021-06-28T17:20:06-0400

(a)

"PE_{top}=mgh=100\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot5\ m=4900\ J."

(b)

"KE_{bottom}=PE_{top}=4900\ J."

(c)

"KE_{bottom}=\dfrac{1}{2}mv^2,""v=\sqrt{\dfrac{2KE_{bottom}}{m}},""v=\sqrt{\dfrac{2\cdot4900\ J}{100\ kg}}=9.9\ \dfrac{m}{s}."